Reduce the quadratic form to a canonical form and find its nature

1) 2xy+2yz+2zx


1
Expert's answer
2021-06-01T05:56:22-0400

Given question is 2xy+2yz+2zx

The general equation "a_1x^2+a_2y^2+a_3z^2+a_4 xy+a_5yz+a_6 zx"

For the given equation, "a_1=a_2=a_3=0"

"a_4=2, a_5=2, a_6=2"

So, matrix "A=\begin{bmatrix}
0 & 1 & 1\\
1 & 0 & 1 \\
1 & 1 & 0 \\
\end{bmatrix}"

After the matrix operation,

Let "\lambda" be the eigen value of the equation, so "A-\lambda I=0"

Given question is 2xy+2yz+2zx

The general equation "a_1x^2+a_2y^2+a_3z^2+a_4 xy+a_5yz+a_6 zx"

For the given equation, "a_1=a_2=a_3=0"

"a_4=2, a_5=2, a_6=2"

So, matrix "A=\begin{bmatrix}
0 & 1 & 1\\
1 & 0 & 1 \\
1 & 1 & 0 \\
\end{bmatrix}"

Let the eigen value for the equation be "\lambda," From the characteristic equation,

"|A-\lambda I|=0"

"\begin{bmatrix}
-\lambda & 1 & 1\\
1 & -\lambda & 1 \\
1 & 1 & -\lambda \\
\end{bmatrix}=0"

After the expansion of the matrix,

"\Rightarrow -\lambda^3+3\lambda -2=0"

"\Rightarrow \lambda^3-3\lambda +2=0"

"(\lambda +1)(\lambda +1)(\lambda -2)=0"

"\lambda =-1, \lambda =-1, \lambda =2"

Now, substituting the eigen values,

"\lambda = 2," "x_1=\begin{bmatrix}
 1\\
1 \\
1 \\
\end{bmatrix}"

"\lambda=-1, x_2=\begin{bmatrix}
 -1\\
1 \\
0 \\
\end{bmatrix}"

"\lambda=-1, x_3=\begin{bmatrix}
 0\\
1 \\
-1 \\
\end{bmatrix}"


Now, "P=(y_1,y_2,y_2)"

Applying the transpose of matrix P,

"P^t=\begin{bmatrix}
 y_1\\
 y_2 \\
y_3
\end{bmatrix}"

Now, multiplying the P , A and transpose matrix of P

"PAP^t=\begin{bmatrix}
y_1 & y_2 & y_3 \\
\end{bmatrix}\begin{bmatrix}
-2 & 1 & 1\\
1 & -1 & 1 \\
1 & 1 & -1 \\
\end{bmatrix}\begin{bmatrix}
 y_1\\
 y_2 \\
y_3
\end{bmatrix}"

"=\begin{bmatrix}
-2y_1 +y_2 + y_3 &y_1-y_2+y_3 & y_1+y_2-y_3
\end{bmatrix}\begin{bmatrix}
 y_1\\
 y_2 \\
y_3
\end{bmatrix}"

"=-2y_1^2+y_1y_2+y_1y_3+y_1y_2-y_2^2+y_2y_3+y_1y_3+y_2y_3-y_3^2"

"=-(y_1^2+y_2^2+y_3^2-2y_1y_2-2y_2y_3-2y_3y_1)"

"=-(y_1-y_2-y_3)^2"

The above equation is in the canonical form.

Rank =3

Index = 1